OpenStudy (tester97):

the function y=-0.296x^2+2.7x models the length x and height y that your sisters rabbit can jump in centimeters. how high does the rabbit jump? what is the length of the jump?

OpenStudy (austinl):

Ok, do you have a graphing calculator?

OpenStudy (tester97):

yup ^_^ who doesnt?

OpenStudy (bibby):

do you have an internet connection? http://www.wolframalpha.com/input/?i=y%3D-1.26%2B2.7x

OpenStudy (austinl):

Wrong equation bud.

OpenStudy (tester97):

bibbeh did u realy just ask me that?

OpenStudy (bibby):

I copied it letter for letter. you have to mess with the parentheses

OpenStudy (austinl):

I am trying to figure out how to do this algebraically again. I know how to do it with calculus easily! And there is an x^2 bibsy.

OpenStudy (bibby):

woah what. I literally copied y=-0.296x^2+2.7x into wolfram. Technology has betrayed me. Go forth austin

OpenStudy (austinl):

Okay, okay I think I got this. We can use the formula, \(\displaystyle x=\frac{-b}{2a}\) When we have, \(y=-ax^2+bx\) Then we can plug it back into the original equation and solve for y. That will be your maximum value. Make sense?

OpenStudy (tester97):

so i have to move the b and a over?

OpenStudy (austinl):

No, they are literally given to you in your initial problem. Look at the general form that I posted and look at your equation. Don't they look similar? a&b are merely coefficients.

OpenStudy (tester97):

sorry this is not making any sense to me whatsoever..its just one of those days

OpenStudy (austinl):

\(y=\color{red}{-0.296}x^2+\color{blue}{2.7}x\) \(y=\color{red}{-a}x^2+\color{blue}{b}x\)

OpenStudy (tester97):

oh yea i understood that part

OpenStudy (austinl):

Okay, now plug that into the little formula that I gave you.

OpenStudy (tester97):

\[x= \frac{ -2.7 }{ -0.296 }\]

OpenStudy (austinl):

x=-b/(2a)

OpenStudy (tester97):

whoops i forgot to multiply the denominator by 2

OpenStudy (tester97):

so it would be \[x=\frac{ -2.7 }{2(-0.296) }\]

OpenStudy (austinl):

Excuse me, I believe I made an error.... Maximum value is, \(\displaystyle max=c-\frac{b^2}{4a}\) Where, \(y=-ax^2+bx+c\) In our case, c=0 Sorry about that :(

OpenStudy (tester97):

so pretty much i would just have to square the b and instead of multiplyin a by 2 it would b 4?

OpenStudy (austinl):

Yeah, sorry again.

OpenStudy (tester97):

its ok :D let me work this out rite quick

OpenStudy (tester97):

-6.15?

OpenStudy (tester97):

oh wait no its not negative so it would just b 6.15?

OpenStudy (austinl):

That is what I get, but remember it is -(-6.15), so 6.15

OpenStudy (austinl):

Beat me to it :P

OpenStudy (austinl):

Then to find the distance, you would just find the x intercepts. The zeros.

OpenStudy (tester97):

how do i do that|?

OpenStudy (austinl):

I believe you could use the quadratic formula. I will brb, try that and I will check back shortly.

OpenStudy (tester97):

ok

OpenStudy (tester97):

wait would i need to just find the vertex?

OpenStudy (tester97):

or did we already do that?

OpenStudy (austinl):

That was how high, we want to know how far now.

OpenStudy (tester97):

oh 9.14 cm?

OpenStudy (austinl):

For some reason I keep getting zero.... But on my calculator is giving me 9.12162

OpenStudy (tester97):

ehh about the same thanks austin ^_^

OpenStudy (austinl):

No problem! Good job btw!

terenzreignz (terenzreignz):

Oh, the repugnant smell of uncertainty... I detect a hint of it here :>

OpenStudy (austinl):

Shhhh @terenzreignz I corrected course......

terenzreignz (terenzreignz):

When did I ever doubt you? :D

OpenStudy (tester97):

tj is that u?

terenzreignz (terenzreignz):

Let me check...

terenzreignz (terenzreignz):

<looks around> Yup, I'm pretty sure my name is TJ, one way or another, since those two letters adorn quite a few select nooks in my room :>

OpenStudy (tester97):

oh tj :P